OpenStudy (anonymous):

Write an equation of a line parallel to y = 2/3 x -5 in slope-intercept form that passes through the point (-6, -1). y= -3x - 1 y= 3x + 6 y= 2/3 x - 6 y= 2/3 x + 3

OpenStudy (shadowlegendx):

Bacon

OpenStudy (adrynicoleb):

Lol. Ok so do you know formula?

OpenStudy (anonymous):

I'm serious D:

OpenStudy (shadowlegendx):

I know how to do it but im too lazy ._.

OpenStudy (adrynicoleb):

*the slope formula

OpenStudy (anonymous):

erm.... y2 - y1 __________________ x2 - x1

OpenStudy (anonymous):

help me out? ;-;

OpenStudy (adrynicoleb):

Yesh. That is all I remember right now because my brain is really tired. I'm sorry.

OpenStudy (adrynicoleb):

You should force shadow to help you. *wink wink*

OpenStudy (anonymous):

SHADOW o-o

OpenStudy (adrynicoleb):

*whispers* Bribe him with bacon, that will have to work.

OpenStudy (marigirl):

@Reneestacha Write an equation of a line parallel to y = 2/3 x -5 in slope-intercept form that passes through the point (-6, -1) If the line is parallel, then it will have the same gradient (slope) as y = 2/3 x -5 Therefore, the gradient (slope) is 2/3 now what we can do is form the equation of the parallel line using gradient m=2/3 and that it passes through (-6, -1) Using y-y1=m(x-x1) I will start you off. y-(-1)=2/3(x-(-6)) would u be able to go on from here?

OpenStudy (shadowlegendx):

Nice marigirl ._.